40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8732     Entry time: Thu Jun 20 09:33:42 2013
Author: Steve 
Type: Update 
Category: General 
Subject: cleanup 

Office work benches were cleaned up yesterday. Anti-image filter boards were moved to north wall of the control room. Koji's pd- electronics box  placed next to water dispenser.

The removed ETMY optical table: TMC 4' x 2' x  4" with Aluminum enclosure was placed on table in the east arm.

Attachment 1: tmc3x2.jpg  148 kB  | Hide | Hide all
tmc3x2.jpg
ELOG V3.1.3-